OpenStudy (chihiroasleaf):

for every integer \(x\) defined a function \(f\) with \(f(x)\) is the number of digits of \(x\). e.g : \(f(125) = 3\) and \(f(2012)=4\). the value of \(f\left( 2^{2012} \right) + f\left(5^{2012}\right) \) is

ganeshie8 (ganeshie8):

How many digits are there in \(\large 10^5\) ?

OpenStudy (chihiroasleaf):

6 digits ?

OpenStudy (chihiroasleaf):

is \( f(2^{2012}) + f(5^{2012}) = f( 2^{2012} \times 5^{2012})\) ? I know the number of digits of the right side, but I don't know whether they're equal

ganeshie8 (ganeshie8):

they are not equal

ganeshie8 (ganeshie8):

use this formula number of digits in \(x\) is given by \(\lfloor \log_{10} x \rfloor + 1\)

ganeshie8 (ganeshie8):

so we have use this formula \[f(x) = \lfloor \log_{10} x \rfloor + 1\]

ganeshie8 (ganeshie8):

actually you might be right! let me think a bit more..

OpenStudy (chihiroasleaf):

could you explain about the formula? I mean, how could the function equal to logarithm ?

ganeshie8 (ganeshie8):

how did you say the number of digits in 10^5 is 6 ?

ganeshie8 (ganeshie8):

you figured out the number of digits as 6 because the "exponent" is 5, so there will be 5 zeroes in the number... plus accounting for 1gives you 6, yes ?

ganeshie8 (ganeshie8):

logarithm = exponent

OpenStudy (chihiroasleaf):

yes.., ahh..., I see...

ganeshie8 (ganeshie8):

\[\large \log_{10} 10^{\color{Red}{5}} = \color{Red}{5}\]

ganeshie8 (ganeshie8):

\[\large \log_{10} 10^{\color{Red}{3}} = \color{Red}{3}\]

ganeshie8 (ganeshie8):

\[\large \log_{10} 10^{\color{Red}{k}} = \color{Red}{k}\]

ganeshie8 (ganeshie8):

adding +1 to \(\large \color{red}{k}\) to get the number of digits

ganeshie8 (ganeshie8):

when \(\large \color{red}{k}\) is not an integer, you just take the integer part \[\large\]

OpenStudy (chihiroasleaf):

so, \[f \left( 2^{2012} \right)=\lfloor \log _{10}2^{2012} \rfloor + 1\] how to calculate \(\lfloor \log _{10}2^{2012} \rfloor\) ?

OpenStudy (chihiroasleaf):

the problem should be solved without calculator :(

ganeshie8 (ganeshie8):

we can find it w/o calculator

ganeshie8 (ganeshie8):

\[\begin{align}f \left( 2^{2012} \right) +\color{blue}{f \left( 5^{2012} \right)} &=\lfloor \log _{10}2^{2012} \rfloor + 1 + \color{blue}{\lfloor \log _{10}5^{2012} \rfloor + 1}\\~\\ &=\lfloor \log _{10}(10/5)^{2012} \rfloor + 1 + \color{blue}{\lfloor \log _{10}5^{2012} \rfloor + 1}\\~\\ &=\lfloor \log _{10}10^{2012} - \log _{10}5^{2012} \rfloor + 1 + \color{blue}{\lfloor \log _{10}5^{2012} \rfloor + 1}\\~\\ &=\lfloor 2012 - \log _{10}5^{2012} \rfloor + 1 + \color{blue}{\lfloor \log _{10}5^{2012} \rfloor + 1}\\~\\ &=2012 + \lfloor - \log _{10}5^{2012} \rfloor + 1 + \color{blue}{\lfloor \log _{10}5^{2012} \rfloor + 1}\\~\\ &=\cdots \end{align}\] see if you can finish it off

ganeshie8 (ganeshie8):

Notice you can pull out integers out of greatest integer function because, for any integer \(n\) and real number \(x\) we have : \[\lfloor n+x \rfloor = n + \lfloor x\rfloor\]

OpenStudy (anonymous):

Here's an attempt at approximating... Hopefully you might recall that \(2^{10}=1024\approx1000=10^3\). So, we have \[10\log_{10}2\approx\log_{10}10^3=3~~\implies~~\log_{10}2\approx0.3\] Multiplying by \(2012\) gives \(\log_{10}2^{2012}\approx 603.6\), and so \(\lfloor\log_{10}2^{2012}\rfloor\approx603\), which means \(2^{2012}\) has somewhere around \(604\) digits (the actual answer is \(606\), but I'm not sure how to fix this approach's accuracy...)

OpenStudy (anonymous):

Might have something to do with the fact that I'm rounding down \(1024\) to \(1000\)

OpenStudy (chihiroasleaf):

thank you very much @ganeshie8 :))